Determine the equilibrium quantity: Qd=25-3P; Qs=10+2P 


1
Expert's answer
2022-03-30T14:29:31-0400

Qd = 25-3P

Qs = 10+2P

At equilibrium Qd= Qs

25-3P = 10 + 2P

25-10 = 2P+3P

15 = 5P

"\frac{15}{5} = \frac{5P}{5}"


P = 3

Qd "= 25-3P"

"=25-3(3)"

"=16"

Therefore the quantity at equilibrium is 16.